# TURB-MHD: an open-access database of forced homogeneous magnetohydrodynamic turbulence

### v1

@@ 1 Numerical simulations
Here $p$ is the pressure divided by the fluid density which is constant, $\nu_{\alpha}$ and $\mu_{\alpha}$ are the hyper-viscosity and hyper-resistivity, and $\alpha$ denotes the power of the Laplacian operator employed in the hyper-dissipation. Standard Laplacian dissipation corresponds to the case $\alpha=1$. The forcing $\boldsymbol{F}$ applied to the system is a drag-free Ornstein–Uhlenbeck process, active in the wavenumber band $k\in[2.5,5.0]$ for each MHD simulation. We consider both standard diffusive ($\alpha=1$) and hyper-diffusive ($\alpha=5$) cases, always with $\nu_{\alpha}=\mu_{\alpha}$ (Borue & Orszag, 1995).
Equations (1)-(4) are stepped forwards in time using a second-order Runge-Kutta scheme in a triply periodic $(2\pi)^{3}$ domain. Both the (hyper)viscous and magnetic (hyper)diffusion terms are treated implicitly using an integrating factor. The spatial discretisation was implemented via the standard pseudospectral method with complete dealiasing by the two-thirds rule (Patterson & Orszag, 1971; Canuto et al., 1988).
Further details and mean values of key observables are summarised in table 1. The hereby presented datasets have been used in Capocci et al. (2025) and Capocci (2024) and they are being analysed for future works (Capocci et al., 2024a, b; Hengster et al., 2024; Hengster, 2025).
id $\frac{B_{0}}{B_{\rm rms}}$ $N$ $\alpha$ $E_{u}$ $E_{b}$ $\nu_{\alpha}$ $\varepsilon_{u}$ $\varepsilon_{b}$ $L_{u}$ $\tau$ Re $\dfrac{k_{\text{max}}}{k_{\text{diss}}^{u}}$ $\dfrac{k_{\text{max}}}{k_{\text{diss}}^{b}}$ $\Delta t/\tau$
@@
[…]
The spatial resolution of the simulations can be quantified by both the grid spacing $\Delta{x}=2\pi/N$ and the (hyper-diffusive) Kolmogorov scales $\eta^{u}_{\alpha}=(\nu_{\alpha}^{3}/\varepsilon_{u})^{1/(6\alpha-2)}$, and $\eta^{b}_{\alpha}=(\mu_{\alpha}^{3}/\varepsilon_{b})^{1/(6\alpha-2)}$, where $\varepsilon_{u}$ and $\varepsilon_{b}$ are the mean kinetic and magnetic energy dissipation rates. For adequate resolution we require $\eta^{u}_{\alpha}/\Delta{x}\gtrsim 1.3$ and $\eta^{b}_{\alpha}/\Delta{x}\gtrsim 1.3$ (e.g., Donzis et al., 2008; Wan et al., 2010).
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Fig. 2 shows the time-averaged omnidirectional kinetic and magnetic spectra for each datasets. Here, the employment of hyperviscosity makes the dissipation range more concentrated in the small scales, leading to the sharp fall-off of the spectra. Although we do not apply a magnetic forcing, the magnetic energy spectra of panels (c) and (d) are characterised by a peak in the forcing wavenumber band because of a spectral coupling due the the non-zero BMF (see e.g. sec. 5.1 of Capocci (2024) for further explanation). Moreover, in the kinetic energy spectrum shown in panel (d), the modes associated with small wavenumbers are significantly more excited compared to those in the other configurations. This is due to a partial inverse cascade of kinetic energy, commensurate with the enhanced two-dimensionalization of the fluctuations, as can be expected in presence of a strong BMF. Although the applied forcing scheme differs, the effects of a strong BMF have been studied, for instance, by Alexakis (2011) and Gallet & Doering (2015).
@@ 1.1 Reynolds number estimate for hyperviscous runs
To calculate an effective Reynolds number for the hyperdissipative datasets we follow the methodology described in Buzzicotti et al. (2018). There, the standard integral-scale Reynolds number, based on standard Laplacian dissipation, $Re={UL_{u}}/{\nu}\propto(L_{u}/\eta_{1})^{4/3}$ (e.g., Batchelor, 1970; Pope, 2000) is replaced with one based on the ratio between the integral scale $L_{u}$ and the effective dissipation range scale $I_{d}$. In particular, we employ
@@ 2.1 Datasets history
Datasets A1 and A3 originate from pre-existing MHD datasets with $N=1024^{3}$ grid-points (Oughton, 2022) employing standard and hyperviscous dissipation laws respectively.
@@ 2.2 Database files
  • During the simulation we dumped the vector potential of both velocity and magnetic field in Fourier space every simulation steps. The instantaneous configuration have been called ”cb” and ”ch” (from Complex B-field and Complex H-field). Each dataset is equipped with a text file called selected_snaps_file.txt that indicates the simulation time of the snapshots sampling.
@@
The SMART-Turb portal also comprises of the following databases: TURB-Hel (Biferale et al., 2024) describing helically forced homogeneous and isotropic turbulence, TURB-Rot (Biferale et al., 2020) for rotating turbulence and TURB-Lagr (Biferale et al., 2024) related to Lagrangian particle under turbulence.

### v2

@@ 1 Numerical simulations
Here $p$ is the pressure divided by the fluid density which is constant, $\nu_{\alpha}$ and $\mu_{\alpha}$ are the hyper-viscosity and hyper-resistivity, and $\alpha$ denotes the power of the Laplacian operator employed in the hyper-dissipation. Standard Laplacian dissipation corresponds to the case $\alpha=1$. The forcing $\boldsymbol{F}$ applied to the system is a drag-free Ornstein–Uhlenbeck process, active in the wavenumber band $k\in[2.5,5.0]$ for each MHD simulation. We consider both standard diffusive ($\alpha=1$) and hyper-diffusive ($\alpha=5$) cases, always with $\nu_{\alpha}=\mu_{\alpha}$ (borue1995forced).
Equations (1)-(4) are stepped forwards in time using a second-order Runge-Kutta scheme in a triply periodic $(2\pi)^{3}$ domain. Both the (hyper)viscous and magnetic (hyper)diffusion terms are treated implicitly using an integrating factor. The spatial discretisation was implemented via the standard pseudospectral method with complete dealiasing by the two-thirds rule (PattersonOrszag71; CanutoEA).
Further details and mean values of key observables are summarised in table 1. The hereby presented datasets have been used in damiano2024a and damiano2024thesis and they are being analysed for future works (damiano2024b; damiano2024c; moritz2024; yasminthesis).
id $\frac{B_{0}}{B_{\rm rms}}$ NN $\alpha$ $E_{u}$ $E_{b}$ $\nu_{\alpha}$ $\varepsilon_{u}$ $\varepsilon_{b}$ $L_{u}$ $\tau$ Re $\dfrac{k_{\text{max}}}{k_{\text{diss}}^{u}}$ $\dfrac{k_{\text{max}}}{k_{\text{diss}}^{b}}$ $\Delta t/\tau$ #
@@
Table 1: Simulation parameters and key observables, where $N$ is the number of collocation points in each coordinate, $\alpha$ is the power of $\nabla^{2}$ used in the hyper-diffusion, $E_{u}$ the mean total kinetic energy, $\nu_{\alpha}$ the kinematic (hyper)viscosity, $\varepsilon_{u}$ and $\varepsilon_{b}$ are the kinetic and Joule energy dissipation rates, $L_{u}=(3\pi/4E_{u})\int_{0}^{k_{\text{max}}}dk\ E_{u}(k)/k$ the integral scale, $\tau=L_{u}/\sqrt{2E_{u}/3}$ the large-scale eddy-turnover time, $B_{\rm rms}=\sqrt{2E_{b}}$ the root-mean-square value of the magnetic field fluctuations, and Re is the (integral scale) Reynolds number. Moreover, $k_{\text{diss}}^{u}$ and $k_{\text{diss}}^{b}$ are the wavenumbers associated with the hyperdiffusive Kolmogorov scales $\eta^{u}_{\alpha}=(\nu_{\alpha}^{3}/\varepsilon_{u})^{1/(6\alpha-2)}$ and $\eta^{b}_{\alpha}=(\mu_{\alpha}^{3}/\varepsilon_{b})^{1/(6\alpha-2)}$ respectively calculated in terms of the viscous and Joule dissipation rates, $k_{\text{max}}$ the largest retained wavenumber component after dealiasing, $\Delta t$ the mean of the snapshots sampling intervals, and # indicates the number of snapshots used in the statistics. The magnetic Prandtl number, $Pm=\nu_{\alpha}/\mu_{\alpha}$ defines the ratio between the hyperviscosity and magnetic hyperdiffusivity, equals unity for each dataset. Regarding C10, the parameters $E_{u}$, $E_{b}$, $L_{u}$, $\tau$ and $Re$ , are related to the last stationarity interval of the run corresponding to $t/\tau\in[95,98]$ (see fig.06).
The spatial resolution of the simulations can be quantified by both the grid spacing $\Delta{x}=2\pi/N$ and the (hyper-diffusive) Kolmogorov scales $\eta^{u}_{\alpha}=(\nu_{\alpha}^{3}/\varepsilon_{u})^{1/(6\alpha-2)}$, and $\eta^{b}_{\alpha}=(\mu_{\alpha}^{3}/\varepsilon_{b})^{1/(6\alpha-2)}$, where $\varepsilon_{u}$ and $\varepsilon_{b}$ are the mean kinetic and magnetic energy dissipation rates. For adequate resolution we require $\eta^{u}_{\alpha}/\Delta{x}\gtrsim 1.3$ and $\eta^{b}_{\alpha}/\Delta{x}\gtrsim 1.3$ (e.g., DonzisEA08; WanEA10-accuracy).
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
Fig. 2 shows the time-averaged omnidirectional kinetic and magnetic spectra for each datasets. Here, the employment of hyperviscosity makes the dissipation range more concentrated in the small scales, leading to the sharp fall-off of the spectra. Although we do not apply a magnetic forcing, the magnetic energy spectra of panels (c) and (d) are characterised by a peak in the forcing wavenumber band because of a spectral coupling due the the non-zero BMF (see e.g. sec. 5.1 of damiano2024thesis for further explanation). Moreover, in the kinetic energy spectrum shown in panel (d), the modes associated with small wavenumbers are significantly more excited compared to those in the other configurations. This is due to a partial inverse cascade of kinetic energy, commensurate with the enhanced two-dimensionalization of the fluctuations, as can be expected in presence of a strong BMF. Although the applied forcing scheme differs, the effects of a strong BMF have been studied, for instance, by alexakis2011 and gallet2015.
@@ 1.1 Reynolds number estimate for hyperviscous runs
To calculate an effective Reynolds number for the hyperdissipative datasets we follow the methodology described in BuzzicottiEA18-xfer. There, the standard integral-scale Reynolds number, based on standard Laplacian dissipation, $Re={UL_{u}}/{\nu}\propto(L_{u}/\eta_{1})^{4/3}$ (e.g., BatchelorTHT; Pope) is replaced with one based on the ratio between the integral scale $L_{u}$ and the effective dissipation range scale $I_{d}$. In particular, we employ
@@ 2.1 Datasets history
Datasets A1 and A3 originate from pre-existing MHD datasets with $N=1024^{3}$ grid-points (sean_priv_comm) employing standard and hyperviscous dissipation laws respectively.
@@ 2.2 Database files
  • During the simulation we dumped the vector potential of both velocity and magnetic field in Fourier space every simulation steps. The instantaneous configuration have been called ”cb” and ”ch” (from Complex B-field and Complex H-field). Each dataset is equipped with a text file called selected_snaps_file.txt that indicates the simulation time of the snapshots sampling. Specifically, for datasets A1–A3–C10 time increases with the snapshot index, whereas for A2–A4–C1 it decreases.
@@
The SMART-Turb portal also comprises of the following databases: TURB-Hel (turb-hel) describing helically forced homogeneous and isotropic turbulence, TURB-Rot (TURB-Rot) for rotating turbulence and TURB-Lagr (turb-hel) related to Lagrangian particle under turbulence.